Contribution Margin Willie Company sells 15,000 units at $21 per unit. Variable costs are $14.28 per unit, and fixed costs are $54,400.

Determine (a) the contribution margin ratio, (b) the unit contribution margin, and (c) income from operations.

a. Contribution margin ratio (Enter as a whole number.) %

b. Unit contribution margin (Round to the nearest cent.) $ per unit

c. Income from operations $
Answer #1

Contribution margin=Sales-Variable cost  

=(21-14.28)=$6.72 per unit

Contribution margin ratio=Contribution margin/Sales

=6.72/21=32%

Total Contribution margin=(6.72*15000)=100800

Less:Fixed cost=54400

Income from operations=46400

Contribution margin ratio 32%
Unit contribution margin $6.72 per unit
Income from operations $46400